I have no idea how to solve this problem because all of the examples I found tell me the volume. Since this doesn't I am completely stuck. How do you solve it without knowing the volume?
Tara plans to take a bath but forgets to close the drain. The bath usually takes 10 minutes to fill and 12 minutes to empty. How long will it be before the bath is filled if Tara continues not to notice that the drain is open?
Please help!

Answer by stanbon(75887)

Tara plans to take a bath but forgets to close the drain. The bath usually takes 10 minutes to fill and 12 minutes to empty. How long will it be before the bath is filled if Tara continues not to notice that the drain is open?
------------------
Filling:
Time = 10 min/job ; Rate = 1/10 job/min.
-------------
Emptying:
Time = 12 min/job ; Rate = 1/12 job/min.
------------
Together rate:
Time = x min/job ; Rate = 1/x job/min
EQUATION:
Fill rate - empty rate = together rate
1/10 - 1/12 = 1/x
2/120 = 1/x
x = 120/2
x = 60 minutes (time to fill the tub)
===================
Cheers,
Stan H.